The mother is eight more than twice the age of her son. Six years ago, she was four more than three times her son's age then. How old are they now.

Answer by ankor@dixie-net.com(22740)

let m = Mother's a age
let s = son's
:
Write an equation for each statement
:
The mother is eight more than twice the age of her son.
m = 2s + 8
Six years ago, she was four more than three times her son's age then.
m - 6 = 3(s - 6) + 4
m - 6 = 3s - 18 + 4
m = 3s - 18 + 6 + 4
m = 3s - 8
replace m with (2s+8) from the 1st statement
2s + 8 = 3s - 8
8 + 8 = 3s - 2s
16 = s
then
m = 2(16) + 8
m = 40
:
How old are they now. Mom's 40, son's 16